Find the equation of the line with m = 3 and passing through point (3, 0).
 
 
a) y = 3x – 9
b) y = 3x – 3
c) y = 3x + 9
d) y = 3x + 3

Answer by math_helper(2461)

 
 
Find the equation of the line with m = 3 and passing through point (3, 0).
 
a) y = 3x – 9
b) y = 3x – 3
c) y = 3x + 9
d) y = 3x + 3
———————————————————————
To solve, compare y = mx+b (m=slope, b=y-intercept) to see which of the choices work. They all have m=3 so then the next question is which one passes through (3,0)?
a) y = 3(3) - 9 = 0 —> this equation, y=3x-9, satisfies both conditions: slope=3 and that the line passes through (3,0).
--
So (a) is the answer.